OpenStudy (anonymous):

Solve 5 – 3(2 – x) = 5 + 3(x – 2).

OpenStudy (aroub):

4-2x=8x-16 4+16=8x+2x 20=10x x=2

OpenStudy (anonymous):

it can't be 2...

jhonyy9 (jhonyy9):

5-6+3x=5+3x-6 - you will get 0=0

OpenStudy (aroub):

ye sry so its no solution/ no equation / no true